OpenStudy (anonymous):

PLZ HELP???!!!!! Carissa deposited $4,232.56 into a savings account with an interest rate of 3.4% compounded daily. About how long will it take for the account to be worth $7,000? 48 years, 8month 15 years, 1 months 19 years, 3 months 14 years, 10 months

OpenStudy (mertsj):

\[A=P(1+\frac{r}{n})^{nt}\]

OpenStudy (mertsj):

Plug into that, and solve for t

OpenStudy (anonymous):

ok let me try and tell me if I am right ok

OpenStudy (anonymous):

so is it \[4,232.56(1+0.034/365)^{365t}\]

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

=7000

OpenStudy (anonymous):

I got 19 years 3 months is that right?

OpenStudy (mertsj):

no

OpenStudy (mertsj):

\[7000=4232.56(1+\frac{.034}{365})^{365t}\] \[\frac{7000}{4232.56}=(1.000093151)^{365t}\] \[\ln 1.653845=365t \ln 1.000093151\]

OpenStudy (mertsj):

Solving that yields t = 14.79 .79(12)=9.57 So 14 years 9.57 months or about 14 years 10 months.

OpenStudy (anonymous):

O ok thanks so much

OpenStudy (mertsj):

yw